A sample of 2-butanol containing 0.375 g in 12.5 mL of solvent was analyzed in a polarimeter using a 10.0 cm polarimeter tube. The sample had an optical rotation , a, of -0.405. )A) What is the specific rotation of 2-butanol? (B) A sample of enantiomer of the compound in A had optical rotation of +0.670 in a 10.0 cm polarimeter tube. What is the concentration of the second sample?
Answer #1

specific rotation= observed rotation/ (path length in dceimeter*concentration in g/ml)

Conncentration = 0.375/12.5 g/ml= 0.03 g/ml

l= 10cm =1 dm

specific rotation= -0.405/(1*0.03)= -13.5 deg.c

for the second case

optical rotation = 0.67 and specific rotation -13.5 deg.c

Enantiomer will have specific rotation of +13.5 deg.c

13.5 =0.67/1*concentration

Concentration (g/l) =0.67/13.5=0.04963 g/ml